a machine gun of mass 12 kg fires bullets of mass 25g at the rate of 4 bullets/sec......with a velocity 500m/s...........what force must be applied to the gun to hold it in position.....??????????????????????PLS EXPLAIN!!!

Grade:9

2 Answers

Nikhil panduranga
12 years ago

First of all you should know that all Physics Qs are based on ideal conditions,so the momentum(linear) always remains conserved....

ie m1v1=m2v2

in this case

GUN                BULLET

m1=12kg        m2=25gm=>0.025kg=>(0.025*4)kg per sec

v1=x(let)m/s  v2=500m/s

since momentum is conserved

12*x=0.025*4*500

x=(2.5*4*5)/12

x=4.16667m/s

F=change in momentum,so

12*4.1667-0="GIVES YOU THE CORRECT ANSWER!!"

PLZ APPROVE THE ANSWE

Rahul Kumar
12 years ago

First of all you should know that all Physics Qs are based on ideal conditions,so the momentum(linear) always remains conserved....

ie m1v1=m2v2

in this case

GUN                BULLET

m1=12kg        m2=25gm=>0.025kg=>(0.025*4)kg per sec

v1=x(let)m/s  v2=500m/s

since momentum is conserved

12*x=0.025*4*500

x=(2.5*4*5)/12

x=4.16667m/s

F=change in momentum,so

12*4.1667-0="GIVES YOU THE CORRECT ANSWER!!"